O Unitek College | Nursing an Portal hosted on web1 serv serv O Week Three Quiz Advertiser S+ 16 Question 17 The Honolulu Advertiser stated that in Honolula there was an average of 661 burglaries per 100,000 households in a are 323 homes. Let r be the number of homes that will be burglarized in a year. Compute the probability for -3. Round given year. In the Kohola Drive neighborhood there Select one: o a. 0.3969 O b. 0,0613 C. 0.2001 O d. o.1890 O e. 0.1743 Previous page Next page O Type here to search
Answer #1

here expected number of burgalries in 323 homes =323*661/100000=2.1

hence from poisson approximation P(r=3)=e-2.1*2.13/3! =0.1890

option D is correct